OpenStudy (anonymous):

What is the arc length of the subtending arc for an angle of 60° on a circle of radius 3?

OpenStudy (anonymous):

180 π π3 π2

OpenStudy (anonymous):

Those are infinity signs

OpenStudy (anonymous):

infinity infinity 3 infinity 2

OpenStudy (anonymous):

arclength=radius x angle

OpenStudy (anonymous):

^ in radians

OpenStudy (jdoe0001):

\(\bf \textit{arc's length}=s=\cfrac{\theta r\pi }{180}\quad \begin{cases} \theta\to \textit{angle in degrees}\\ r\to radius \end{cases} \)

OpenStudy (anonymous):

a?

OpenStudy (anonymous):

You tell us.

OpenStudy (anonymous):

You said radius times angle so is it a?

OpenStudy (anonymous):

jdoe is it infi 3?

OpenStudy (jdoe0001):

@DaWizjr that is correct, FOR a radian angle in this case, you're using 60 degrees thus

OpenStudy (anonymous):

so its infi 3?

OpenStudy (jdoe0001):

infi 3? hmm what does \(\bf \textit{arc's length}=s=\cfrac{\theta r\pi }{180}\quad \begin{cases} \theta\to \textit{angle in degrees}\\ r\to radius \end{cases} \) give you anyway?

OpenStudy (anonymous):

60rinfi/180=

OpenStudy (jdoe0001):

recall that the angle is 60, and the radius is 3 thus \(\bf \textit{arc's length}=s=\cfrac{\theta r\pi }{180}\quad \begin{cases} \theta\to \textit{angle in degrees}\\ r\to radius \end{cases}\qquad s=\cfrac{60\cdot 3\cdot \pi }{180}\)

OpenStudy (anonymous):

I see now

OpenStudy (anonymous):

so its infi

OpenStudy (jdoe0001):

infi? infinity?

OpenStudy (anonymous):

didnt multiply the radius

OpenStudy (anonymous):

yea infinity

OpenStudy (jdoe0001):

let us think about that we have a circle, a radius of 3, an angle of 60 degrees 

OpenStudy (anonymous):

so its not infinite

OpenStudy (jdoe0001):

nope..... so.... \(\bf \begin{cases} \theta\to \textit{angle in degrees}\\ r\to radius \end{cases}\qquad s=\cfrac{60\cdot 3\cdot \pi }{180}\) is?

OpenStudy (anonymous):

I got just 1 infinity

OpenStudy (jdoe0001):

1 infinity? as opposed to 5 infinities? thought there was only one... well, there's a positive and a negative in cartesian terms anyway so... think about it... does that arc look like it goes on and has no end?

OpenStudy (jdoe0001):

unless you're calling the \(\huge \pi \) infinity, is called \(\huge PI\) as opposed to "pie" btw

OpenStudy (anonymous):

no maybe im looking at it wrong

OpenStudy (jdoe0001):

\(\huge \infty\) is infinity, \(\huge \pi\) is "pi"

OpenStudy (anonymous):

ok

OpenStudy (jdoe0001):

\(\bf \textit{arc's length}=s=\cfrac{\theta r\pi }{180}\quad \begin{cases} \theta\to \textit{angle in degrees}\\ r\to radius \end{cases}\qquad s=\cfrac{60\cdot 3\cdot \pi }{180} \\ \quad \\ s=\cfrac{\cancel{180}\pi }{\cancel{180}}\implies s=\pi\implies s\approx 3.1416\)

OpenStudy (anonymous):

yea I know pi is 3.14

OpenStudy (jdoe0001):

there are even a few movies on \(\huge \pi\) btw you may like them, is a bit longer than 3.1416 but rounded up is about that much, or 3.14

OpenStudy (anonymous):

and then I divide and still get 3.14

OpenStudy (anonymous):

but pi is infinity

OpenStudy (jdoe0001):

hmmm \(\huge \pi\) is an irrational number, is not 3, is not 4 is more than 3 is less than 4 but is not infinite

OpenStudy (anonymous):

so the answer is not any infinities?

OpenStudy (jdoe0001):

hmmm have you covered linear simplifications yet?

OpenStudy (anonymous):

Nah I haven't

OpenStudy (jdoe0001):

ahh.. .well... then you may want to start there I'd think this exercise doesn't apply to you then

OpenStudy (anonymous):

Im sorry I wasnt paying attention

OpenStudy (anonymous):

the answer is 3.14

OpenStudy (anonymous):

pi

OpenStudy (anonymous):

@jdoe0001

OpenStudy (anonymous):

my fault for getting the pi symbol and infiniti symbol confused

OpenStudy (anonymous):

@jdoe0001